OpenStudy (anonymous):

1 + sec^2(x) +sin ^2 (x) = sec^2(x)

OpenStudy (psymon):

I didnt instantly see it, so looking at it, lol.

OpenStudy (psymon):

Are you sure that's correct?

OpenStudy (psymon):

Whoops, compsomehow crashed. So is this an identity or solving for x?

OpenStudy (anonymous):

Its 1 +sec ^ 2 (x) sin^2(x) = sec^2(x) and its proving trig equations

OpenStudy (psymon):

Oh! Its times? You put plus o.o

OpenStudy (anonymous):

Sorry

OpenStudy (psymon):

Alrighty, I'll look at it again then :P

OpenStudy (psymon):

Oh, this is pretty easy then. Change sec^2(x) into it's inverse. if you know what I mean.

OpenStudy (anonymous):

Yeah 1/cos

OpenStudy (psymon):

Right. So this would give you: \[1+\frac{ 1 }{ \cos ^{2}x }*\sin ^{2}x = \sec ^{2}x\] Think you can see it from there?

OpenStudy (anonymous):

Yeah thats what I have written down but from there I dont know how to solve

OpenStudy (psymon):

Well, what is sin/cos?

OpenStudy (anonymous):

tan

OpenStudy (psymon):

Right. So you have 1 + tan^2(x), which is sec^2(x)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

If thats it thanks and I became your fan for being helpful.

OpenStudy (psymon):

Glad to help :3

OpenStudy (anonymous):

can I ask you to help on one more question ?